# Snail: Secure Single Iteration Localization

### v1

@@ 6. Single Iteration Localization
The localization algorithms considered (i.e. Algorithm 2) are iterative by nature; they run a sequence of gradient descent steps until convergence is reached. Furthermore, each iteration requires computing the (pseudo) inverse of a matrix which is itself an iterative process when using the popular algorithm from (Demmel and Kahan, 1990), used in the Eigen and LAPACK libraries to reduce the bidiagonal form to singular values. Executing such iterative algorithms in a data-oblivious setting (like under MPC) have a critical efficiency failure: a fixed number of iterations must be executed regardless of the rate of convergence. Either work is wasted by running more than necessary iterations, or completeness of the solution is degraded by running too few.
Our approach maintains the invariant that control flow does not depend on secret data but avoids wasting work by running on multiple inputs. The outer iterative algorithm, LM optimization via gradient decent, is broken apart such that each step runs independently. Then, stringing the steps from multiple localizations together hides the convergence rate of each individual localization. The inner algorithm, the SVD, is addressed through finding that in practice there is a constant and optimal of number of iterations which does not depend on secret data. We call the combination of the two techniques Single Iteration Localization and follow with details.
@@
There is still one remaining problem in SIL as described. A single gradient descent step requires inverting a matrix which is itself an iterative algorithm whose convergence is data-dependent. Just as before, this suggests an upper bound of iterations is required however the previous solution to reduce the number of iterations does not work as this problem is nested within the previous. It turns out there is an optimal number of iterations which is constant and known beforehand, and thus is easy to encode in the garbled circuit. To see why, we describe more about how the SVD of a matrix is computed, the underlying operation in matrix inversion. To compute the SVD of a matrix it is first reduced to an upper bidiagonal form using Householder transformations. Then, the SVD is computed from the bidiagonal matrix using QR factorization, eliminating the upper diagonal entries as it sweeps across the diagonal in an iterative fashion. The number of iterations taken to reduce the bidiagonal form to singular values depends on the values of the matrix being decomposed. A general rule of thumb is two QR sweeps per singular value (Parlett, 1998, p. 165,) but a bidiagonal matrix whose diagonal entries are equal to the superdiagonal takes fewer QR factorization iterations (Demmel and Kahan, 1990, Table 2.) The key insight we rely on is that the number of singular values of the matrix is always six as each corresponds to a physical degree of freedom. The question is then if localization algorithms require inverting a matrix which, after householder transformations (the first step in computing the SVD), result in diagonal entries equal to the superdiagonal. The geometric implication of this phenomenon with respect to localization is not immediately clear, however, for every experiment performed including real world images and those from the ETH3D dataset, there are no cases where a single diagonal entry is within ±$\pm 10^{-5}$ of the superdiagonal, much less all diagonal entries. This suggests the number of iterations required to compute the SVD is a constant known ahead of time. Namely, twelve QR sweeps are required for convergence, two for each singular value, and there are six singular values each corresponding to a physical degree of freedom. The takeaway is that the fixed upper bound of iterations required to compute the SVD can be set as the optimal number of iterations which in practice does not depend on the secret inputs.
@@
As we later quantify, Single Iteration Localization is faster than executing one large garbled circuit for each localization run when considering a stream of images, however it comes at the cost of additional round complexity. Instead of participants interacting once per image to share a garbled circuit, they must instead do so for every intermediate gradient decent step. In practice round complexity is not an issue as connectivity is already assumed. Furthermore, the network RTT time is much faster than the time to run a localization iteration, so the effect of the additional rounds is minimal.
@@ 6.1. Implementation
We compare the baseline against Single Iteration Localization, also implemented for GM and LM in both ABY and EMP in C++, all of which are available under the MIT license [^2] https://github.com/secret-snail/localization-server. The ABY implementation is 3,887 lines of code, EMP is 1,924 lines, and reference plaintext implementation used for automated testing is 848 lines. Each implementation is an independent library that includes all necessary linear algebra operations depending only on the respective MPC library, including MPC-based adaptations of the SVD algorithm from (Demmel and Kahan, 1990).
@@ 7.1. Security Analysis of Single Iteration Localization
Recall the key insight of Single Iteration Localization is to break each iteration of gradient descent apart. To estimate the pose of a single image, SIL must be invoked repeatedly, until the pose converges. If computing the pose of one image, this would still reveal the number of invocations to the MPC participants, which is a security problem as previously discussed. But if computing the pose on a stream of images, the MPC participants cannot tell which pose refinement iterations belong to which images. This hides how many iterations were required to converge for each image. If the MPC participants also do not know how many images were used as input, they cannot infer the number of iterations per image, which addresses the leakage. Decoupling the number of iterations from the higher level localization executions is the key to the security of SIL.
There is one issue with the security argument as described. We have not quantified how many images count as a “stream” or more specifically, how many times SIL must be invoked to hide the association of gradient decent steps to input images. For example, a stream containing one single image which requires three gradient decent iterations to compute the pose doesn’t offer very good security properties because the MPC participants can see only three iterations were performed and so the initial pose estimate must have been very close to the true pose. It turns out, it depends on a public parameter of the LM and GN localization algorithms, namely the upper bound on the number of gradient decent steps. OpenCV’s implementation of LM localization defines a maximum number of iterations of 20 meaning if the pose hasn’t converged after 20 iterations, localization will halt. It follows that if at least 20 iterations are performed, the probability of an adversary guessing the correct number of iterations per image is no better than randomly guessing.
More generally, if SIL is invoked $o$ times and the publicly known constant upper bound of iterations is $c$, the number of images which could have been used as input is between $\frac{o}{c}$ and $o$. Thus, the probability of an adversarial offload server $A$ correctly guessing the number of input images $i$ from the number of iterations it ran, $o$, is given by:
@@
where $p(\kappa)$ is a polynomial in security parameter $\kappa$, an artifact which appears due to the negligible advantage an attacker has when distinguishing garbled circuit wire labels. Thus, we conclude that Single Iteration Localization is secure when called a minimum of $c$ times. In practice, the default maximum from OpenCV is $c=20$ and since localization is often called repeatedly (more than 20 times), this is not a difficult requirement to meet.
@@
In summary, Single Iteration Localization does not reveal any information about the input features (both image and map), or the pose when run more than $c$ times in series which in practice is 20. This is true for any input including the case when the same input image and map is repeating (corresponding to a stationary camera) or for inputs that change over time (a moving camera). Previous attacks on privacy preserving localization rely on exploiting knowledge of the feature locations to reconstruct the input image. Such attacks are not possible on Single Iteration Localization as the image features, map features, and all other partial information remains confidential. Next we discuss performance implications of our approach.
@@ 7.2. Performance Evaluation
Server side performance is evaluated using a desktop with an 11th generation Intel® core-series CPU. Each MPC participant is a process on the test machine communicating over localhost where bandwidth is limited to 2.5 Gbps and latency is introduced artificially where noted using traffic control. The rationale for 2.5 Gbps being it is the maximum theoretical bandwidth of WiFi 6E, the latest standard at the time of writing. Localization is performed on feature locations from the ETH3D dataset (Schops et al., 2017) where each measurement reported is an average of randomly selected points where the same random sequence is used across measurements. Measurements for which convergence was not achieved are not included in the results (in these cases the plaintext implementation also did not converge). All times reported are wall clock. Implementations are not multithreaded due to their high bandwidth requirements as later shown.
@@
Next, we consider network communication between client and offload servers for the client-server computation offload setting. EMP, being based on the half-gates garbled circuits protocol (Zahur et al., 2015), shares secret inputs via a correlated oblivious transfer protocol. Localization is a unique case where neither generator nor evaluator (the two roles in the protocol the offload servers play) have the secret input data as they are computing on behalf of a client. The natural approach to this situation is the generator sends all wire label pairs to the client, who then forwards the appropriate label from the pair, based on their secret bits, to the evaluator. This requires the client receive $2\kappa$ bits per input bit and send $\kappa$ bits per input bit where $\kappa$ is a security parameter, the size of the wire labels (128 bits by default in EMP’s semi-honest protocol). In the case of six input point correspondences and EMP’s default security parameters, this amounts to the client sending 123 Kb and receiving 246 Kb. We improve upon the natural approach by noticing if the client knows the global circuit delta (a part of the free XOR optimization (Kolesnikov and Schneider, 2008)), and circuit seed, they may generate labels for the evaluator themselves. From a security perspective this is not an issue, as the client is considered a trusted third party and is allowed to learn all inputs and outputs of the computation. The seed based approach reduces data the client receives to a constant $2\kappa$ bits (circuit delta and seed) while data sent remains $\kappa$ bits per input bit (wire labels). In the case of six input point correspondences this amounts to 123 Kb of total communication instead of 369 Kb (123 + 246) via the natural approach.
7.3. Scalability Limitations
@@
Thus far, performance has been evaluated on small input sizes, between six and twelve input point correspondences. While this range is representative of small input sizes for the localization methods considered and is used by other work (Schneider et al., 2018), the quality of localization can increase with larger input sizes notably when input data is noisy. Feature detection algorithms like SIFT (Lowe, 2004) typically extract many more features, around 1000 points per image, but the number which can be matched to 3D map features is typically lower. Thus, we consider the performance of input sizes up to 256 points which is reasonable for popular feature detection algorithms. Prior localization work considers between 100 and 300 points (Qin et al., 2018). Figure 6(a) focuses on the interaction between the two offload servers, namely how latency between the two affects runtime and the relationship between input size and network communication between offload servers.
@@
Comparing the scalability of GN to LM, both algorithms invert a matrix on every optimization iteration; the key difference is the size of the matrix each inverts. GN computed the pseudo-inverse of a $2n\times 6$ Jacobian matrix $\mathbf{J}$ where $n$ is the number of input point correspondences and each row represents the partial derivative with respect to one degree of freedom in the pose. LM instead inverts a $6\times 6$ matrix ($\mathbf{J}^{\top}\mathbf{J}$ plus an offset along the diagonal from (Fletcher, 1971)). Because of the performance impact of inverting large matrices, LM is expected to perform better than GN on large inputs. This behavior is not explicitly clear for small numbers of points shown in Figure 6, but becomes important as the input size scales as shown in Figure 6(a). Because LM also outperforms GN on small input sizes, LM is in general the better approach. We see each LM iteration of SIL takes roughly 20 seconds to compute on 256 input points in Figure 6(a). Assuming 8 iterations to converge (a realistic value observed in the ETH3D dataset), localization completes in 160 seconds.
@@
The performance envelope and high network utilization of the proposed methods are such that localization is best performed using a small number of high quality feature correspondences. As such, applying these methods in practice likely requires input pre-processing to eliminate outlier or even unnecessary inlier point correspondences.
Before continuing, we briefly mention performance related to trusted execution environments (TEEs) like Intel® SGX. Memory requirements to perform localization on the number of points we consider is small enough to fit into even first generation secure enclaves, thus we expect performance in TEEs to be near native (plaintext) speed, making Single Iteration Localization orders of magnitude slower than the 10s of milliseconds required to compute in the clear. The “sluggish” performance of MPC compared to TEE reflects the cost of eliminating side channels, reducing requirements for specialized hardware, and eliminating trust required of hardware vendors.
@@ 7.4. Proof of Concept: Turbo the Snail
To demonstrate the feasibility of privacy preserving localization, we built a Raspberry Pi-based robot acting as a lightweight client which offloads localization. The robot is equipped with one RGB camera and is programmed to move to a target location defined within its environment. The robot uses the AprilTag marker library (Olson, 2011) to detect marker images printed on paper in its environment. The marker corners are detected in the camera image and then matched to a ground truth 3D map of the marker’s position, yielding the 2D and 3D features used in localization. Once the pose is known, the robot moves to a predefined target position using position-based visual servoing (Chaumette and Hutchinson, 2006) via the ViSP library (Marchand et al., 2005).

### v2

@@ 6. Single Iteration Localization
The localization algorithms considered (i.e. Algorithm 2) are iterative by nature; they run a sequence of gradient descent steps until convergence is reached. Furthermore, each iteration requires computing the (pseudo) inverse of a matrix which is itself an iterative process when using the popular algorithm from Demmel and Kahan (Demmel and Kahan, 1990), used in the Eigen and LAPACK libraries to reduce the bidiagonal form to singular values. Executing such iterative algorithms in a data-oblivious setting (like under MPC) have a critical drawback: a fixed number of iterations must be executed regardless of the rate of convergence. Either work is wasted by running more than necessary iterations, or completeness of the solution is degraded by running too few.
Our approach maintains the invariant that control flow does not depend on secret data but avoids wasting work by running on multiple inputs. The outer iterative algorithm, LM optimization via gradient decent, is broken apart such that each step runs independently. Then, stringing the steps from multiple localizations together hides the convergence rate of each individual localization. The inner algorithm, the SVD, is addressed through finding that in practice there is a constant and optimal of number of iterations which does not depend on secret data. We call the combination of the two techniques Single Iteration Localization.
@@
There is still one remaining problem in SIL as described. A single gradient descent step requires inverting a matrix which is itself an iterative algorithm whose convergence is data-dependent. Just as before, this suggests an upper bound of iterations is required however the previous solution to reduce the number of iterations does not work as this problem is nested within the previous. It turns out there is an optimal number of iterations which is constant and known beforehand, and thus is easy to encode in the garbled circuit. To see why, we describe more about how the SVD of a matrix is computed, the underlying operation in matrix inversion. To compute the SVD of a matrix it is first reduced to an upper bidiagonal form using Householder transformations. Then, the SVD is computed from the bidiagonal matrix using QR factorization, eliminating the upper diagonal entries as it sweeps across the diagonal in an iterative fashion. The number of iterations taken to reduce the bidiagonal form to singular values depends on the values of the matrix being decomposed. A general rule of thumb is two QR sweeps per singular value (Parlett, 1998, p. 165,) but a bidiagonal matrix whose diagonal entries are equal to the superdiagonal takes fewer QR factorization iterations (Demmel and Kahan, 1990, Table 2.) The key insight we rely on is that the number of singular values of the matrix is always six as each corresponds to a physical degree of freedom. The question is then if localization algorithms require inverting a matrix which, after householder transformations (the first step in computing the SVD), result in diagonal entries equal to the superdiagonal. The geometric implication of this phenomenon with respect to localization is not immediately clear, however, for every experiment performed including real world images and those from the ETH3D dataset, there are no cases where a single diagonal entry is within ±$\pm 10^{-5}$ of the superdiagonal, much less all diagonal entries. If there is a case where more QR sweeps are required, the accuracy of the localization may be affected, however, this suggests the number of iterations required to compute the SVD is a constant known ahead of time. Namely, twelve QR sweeps are required for convergence, two for each singular value, and there are six singular values each corresponding to a physical degree of freedom. The takeaway is that the fixed upper bound of iterations required to compute the SVD can be set as the optimal number of iterations which in practice does not depend on the secret inputs.
@@
As we later quantify, Single Iteration Localization is faster than executing one large garbled circuit for each localization run when considering a stream of images, however it comes at the cost of additional round complexity. Instead of participants interacting once per image to share a garbled circuit, they must instead do so for every intermediate gradient decent step. In practice round complexity is not an issue as connectivity is already assumed. Furthermore, the network RTT time is much faster than the time to run one iteration, so the effect of the additional rounds is minimal.
@@ 6.1. Implementation
We compare the baseline against Single Iteration Localization, also implemented for GN and LM in both ABY and EMP in C++, all of which are available under the MIT license [^2] https://github.com/secret-snail/localization-server. The ABY implementation is 3,887 lines of code, EMP is 1,924 lines, and reference plaintext implementation used for automated testing is 848 lines. Each implementation is an independent library that includes all necessary linear algebra operations depending only on the respective MPC library, including MPC-based adaptations of the SVD algorithm from Demmel and Kahan (Demmel and Kahan, 1990).
@@ 7.1. Security Analysis of Single Iteration Localization
Recall the key insight of Single Iteration Localization is to break each iteration of gradient descent apart. To estimate the pose of a single image, SIL must be invoked repeatedly, until the pose converges. If computing the pose of one image, this would still reveal the number of invocations to the MPC participants, which is a security problem as previously discussed. But if computing the pose on a stream of images, the MPC participants cannot tell which pose refinement iterations belong to which images. This hides how many iterations were required to converge for each image. If the MPC participants also do not know how many images were used as input, they cannot infer the number of iterations per image, which addresses the leakage. Decoupling the number of iterations from the higher level localization executions is the key to security.
There is one issue with the security argument as described. We have not quantified how many images count as a “stream” or more specifically, how many times SIL must be invoked to hide the association of gradient decent steps to input images. For example, a stream containing one single image which requires three gradient decent iterations to compute the pose doesn’t offer very good security properties because the MPC participants can see only three iterations were performed and so the initial pose estimate must have been very close to the true pose. It turns out, it depends on a public parameter of the LM and GN localization algorithms, namely the upper bound on the number of gradient decent steps. OpenCV’s implementation of LM localization defines a maximum number of iterations of 20 meaning if the pose hasn’t converged after 20 iterations, localization will halt. It follows that if at least 20 iterations are performed, the probability of an adversary outputting the correct number of iterations per image is no better than guessing.
More generally, if SIL is invoked $o$ times and the publicly known upper bound of iterations is $c$, the number of images which could have been used as input is between $\frac{o}{c}$ and $o$ (where $o>c$). Thus, the probability of an adversarial offload server $A$ correctly outputting the number of input images $i$ from the number of iterations it ran, $o$, is given by:
@@
where $p(\kappa)$ is a polynomial in security parameter $\kappa$, an artifact which appears due to the negligible advantage an attacker has when distinguishing garbled circuit wire labels. Thus, we conclude that Single Iteration Localization is secure when invoked a minimum of $c$ times. In practice, the default maximum from OpenCV is $c=20$ and since localization is often called repeatedly (more than 20 times) on camera frames, this is not a difficult requirement to meet.
@@
In summary, Single Iteration Localization does not reveal any information about the input features (both image and map), or the pose when run more than $c$ times in series which in practice is 20. This is true for any input including the case when the same input image and map is repeating (corresponding to a stationary camera) or for inputs that change over time (a moving camera). Previous attacks on privacy preserving localization rely on exploiting knowledge of the feature locations to reconstruct the input image. Such attacks are not possible on Single Iteration Localization as the image features, map features, and all other partial information remains confidential.
@@ 7.2. Performance Evaluation
Server side performance is evaluated using a desktop with an 11th generation Intel® core-series CPU. Each MPC participant is a process on the test machine communicating over localhost where bandwidth is limited to 2.5 Gbps and latency is introduced artificially where noted using traffic control. The rationale for 2.5 Gbps being it is the maximum theoretical bandwidth of WiFi 6E, the latest standard at the time of writing. Localization is performed on features from the ETH3D dataset (Schops et al., 2017) where each measurement is an average of randomly selected points where the same random sequence is used across measurements. Measurements for which convergence was not achieved are not included in the results (in these cases the plaintext implementation also did not converge). Reported times are wall clock and implementations are not multithreaded due to their high bandwidth requirements as later shown.
@@
Next, we consider network communication between client and offload servers for the client-server computation offload setting. EMP, being based on the half-gates garbled circuits protocol (Zahur et al., 2015), shares secret inputs via a correlated oblivious transfer protocol. Localization is a unique case where neither generator nor evaluator (the two roles in the protocol the offload servers play) have the secret input data as they are computing on behalf of a client. The natural approach to garbled circuits is the generator sends all wire label pairs to the client, who then forwards the appropriate label from the pair, based on their secret bits, to the evaluator. This requires the client receive $2\kappa$ bits per input bit and send $\kappa$ bits per input bit where $\kappa$ is a security parameter, the size of the wire labels (128 bits by default in EMP’s semi-honest protocol). In the case of six input point correspondences and EMP’s default security parameters, this amounts to the client sending 123 Kb and receiving 246 Kb. We improve upon this by noticing if the client knows the global circuit delta (a part of the free XOR optimization (Kolesnikov and Schneider, 2008)), and circuit seed, they may generate labels for the evaluator themselves. From a security perspective this is not an issue, as the client is considered a trusted third party and is allowed to learn all inputs and outputs of the computation in the offload setting we consider. The seed based approach reduces data the client receives to a constant $2\kappa$ bits (circuit delta and seed) while data sent remains $\kappa$ bits per input bit (wire labels). In the case of six input point correspondences this amounts to 123 Kb of total communication instead of 369 Kb (123 + 246) via the natural approach.
7.3. Limitations
@@
Thus far, performance has been evaluated on small input sizes, between six and twelve input point correspondences. While this range is representative of small input sizes for the localization methods considered and is used by other work (Schneider et al., 2018), the quality of localization can increase with larger input sizes notably when input data is noisy. Feature detection algorithms like SIFT (Lowe, 2004) typically extract many more features, around 1000 points per image, but the number which can be matched to 3D map features is typically lower. Thus, we consider the performance of input sizes up to 256 points which is reasonable for popular feature detection algorithms. Prior localization work considers between 100 and 300 points (Qin et al., 2018). Figure 7(a) focuses on the interaction between the two offload servers, namely how latency affects runtime and the relationship between input size and network communication.
Localization even on small input sizes takes around 10 seconds to compute, which likely precludes time-sensitive applications like AR/VR because the displacement distance, or error accumulated in the pose as estimated by less accurate means, grows beyond what is acceptable before the more accurate pose may be computed. This makes Single Iteration Localization better suited for less time sensitive applications like robotics or offline processing, for example those described in Section 2.3.
@@
Comparing the scalability of GN to LM, both algorithms invert a matrix on every optimization iteration; the key difference is the size of the matrix each inverts. GN computes the pseudo-inverse of a $2n\times 6$ Jacobian matrix $\mathbf{J}$ where $n$ is the number of input point correspondences and each row represents the partial derivative with respect to one degree of freedom in the pose. LM instead inverts a $6\times 6$ matrix ($\mathbf{J}^{\top}\mathbf{J}$ plus an offset along the diagonal from (Fletcher, 1971)). Because of the performance impact of inverting large matrices, LM is expected to perform better than GN on large inputs. This behavior is not explicitly clear for small numbers of points shown in Figure 6, but becomes important as the input size scales as shown in Figure 7(a). Because LM also outperforms GN on small input sizes, LM is in general the better approach. We see each LM iteration of SIL takes roughly 20 seconds to compute on 256 input points in Figure 7(a). Assuming 8 iterations to converge (a realistic value observed in the ETH3D dataset), localization completes in 160 seconds.
@@
In Single Iteration Localization, we propose executing each iteration of the outer gradient descent algorithm independently. We argue this approach is secure if invoked a minimum number of times, a number which depends on a public parameter of the plaintext algorithm, as discussed in Section 7.1. This is amenable to repeated invocation however it could be seen as a limitation of in the case where a single or few input images must be localized using a large stream size. In this case, however, SIL is no slower than the naïve data-oblivious adaptation and still benefits from running fewer SVD iterations driven by the analysis in Section 6.
The performance envelope and high network requirements of the proposed methods are such that localization is best performed using a small number of high quality feature correspondences. As such, applying these methods likely requires input pre-processing to eliminate outlier or even unnecessary inlier point correspondences.
Before continuing, we briefly mention performance related to trusted execution environments (TEEs) like Intel® SGX. Memory requirements to perform localization on the number of points considered is small enough to fit into secure enclaves, thus we expect performance in TEEs to be near native (plaintext) speed, making Single Iteration Localization orders of magnitude slower than the 10s of milliseconds required to compute in the clear. The “sluggish” performance of MPC compared to TEE reflects the cost of eliminating side channels, reducing requirements for specialized hardware, and eliminating trust required of hardware vendors.
@@ 7.4. Proof of Concept: Turbo the Snail
To demonstrate the feasibility of privacy preserving localization, we built a Raspberry Pi-based robot acting as a lightweight client which offloads localization. The robot is equipped with one RGB camera and uses its onboard WiFi module to communicate with the offload servers, a maximum bandwidth of 100 Mbps. The robot is programmed to move to a target location defined within its environment, leveraging the AprilTag library (Olson, 2011) to detect marker images printed on paper in its environment. The marker corners are detected in the camera image and then matched to a ground truth 3D map of the marker’s position, yielding the 2D and 3D features used in localization. Once the pose is known, the robot moves to a predefined target position using position-based visual servoing (Chaumette and Hutchinson, 2006) via the ViSP library (Marchand et al., 2005).
@@ 8. Conclusion
Acknowledgements.
Thank you anonymous shepherd and reviewers. This work has been supported by NSF awards CCF-2217070 and CNS-1909769.